Complete solution of d^2y/dx^2 +2dy/dx +y=sin 2x



1
Expert's answer
2021-05-07T14:35:57-0400

We have given the differential equation,


"\dfrac{d^2y}{dx^2} +2\dfrac{dy}{dx} +y=sin 2x"


"(D^2+2D+1)y = sin2x"


Auxiliary equation is:

"m^2+2m+1 = 0"

"(m+1)^2 = 0"

"m = -1,-1"

"CF = (C_1+C_2x)e^{-x}"


"PI = \dfrac{sin2x}{D^2+2D+1}" "= sin2x\dfrac{1}{-4+2D+1}"


"= sin2x\dfrac{2D-3}{4D^2-9} = -\dfrac{1}{25}(2D-3)sin2x = -\dfrac{1}{25}(4cos2x-3sin2x)"


Hence the solution of given differential equation is :


"y = (C_1+C_2x)e^{-x}-\dfrac{1}{25}(4cos2x-3sin2x)"